Solve. –4(x + 10) + 2x = 3x + 0
Answer by jai_kos(139)

Solution:
Given – 4(x + 10) + 2 x = 3x + 0
 - 4x – 40 + 2x = 3x + 0
Get the entire “x” terms to one side and numbers to the other side.
-4x + 2x – 3x = 40
- 5x = 40
Now divide the above equation by (-5), we get
x = (40) / -5
x = -8
When solving the above equation, we get x = 8